Calculate the cell potential of the following cell at 298K.
Zn(s)|Zn^2+(0.6 M)||Cu^2+(0.3 M)|Cu(s)
E0 (Zn^2+| Zn) = -0.76 volt, E0(Cu^2+ |Cu) =0.34 volt.
1
Expert's answer
2012-09-07T09:09:02-0400
Cell notation = anode||cathode (oxidation||reduction)

Zn(s) ------> Zn2+ + 2e- Eo = -0.76 V
Cu2+ + 2e- --------> Cu Eo = 0.34 V

Eo = 0.76 + 0.34 = 1.1 V

E = Eo - 0.0591/2 * log ([Zn2+]/[Cu2+])

E = 1.1 - 0.0296 log ([0.6] / [0.3])

E = 1.1 - 0.00891 = 1.091 V